OpenStudy (anonymous):

Rationalize the denominator of sqrt -9 / (4-7i) - (6-6i)

OpenStudy (anonymous):

simplify numerator and denominator and multiply by complex conjugate of denominator\[=\frac{\sqrt{-9}}{4-7i-6+6i}=\frac{3i}{-2-i}*\frac{-2+i}{-2+i}\](continued)

OpenStudy (anonymous):

\[=\frac{-6i+3i^2}{4+1}=\frac{-3-6i}{5}=-\frac{3}{5}-\frac{6i}{5}\]